# Calibrating Verbalized Confidence with Self-Generated Distractors

### v1

@@ Abstract
Calibrated confidence estimates are necessary for large language model (LLM) outputs to be trusted by human users. While LLMs can express their confidence in human-interpretable ways, verbalized LLM-generated confidence scores have empirically been found to be miscalibrated, reporting high confidence on instances with low accuracy and thereby harming trust and safety. We hypothesize that this overconfidence often stems from a given LLM’s heightened suggestibility when faced with claims that it encodes little information about; we empirically validate this hypothesis, finding more suggestibility on lower-accuracy claims. Building on this finding, we introduce Distractor-Normalized Coherence (DiNCo), which estimates and accounts for an LLM’s suggestibility bias by having the model verbalize its confidence independently across several self-generated distractors (i.e. alternative claims), and normalizes by the total verbalized confidence. To further improve calibration, we leverage generator-validator disagreement, augmenting normalized validator confidence with a consistency-based estimate of generator confidence. Here, we frame the popular approach of self-consistency as leveraging coherence across sampled generations, and normalized verbalized confidence as leveraging coherence across validations on incompatible claims, allowing us to integrate these complementary dimensions of coherence into DiNCo. Moreover, our analysis shows that DiNCo provides less saturated – and therefore more usable – confidence estimates, and that further sampling alone cannot close the gap between DiNCo and baselines, with DiNCo at 10 inference calls outperforming self-consistency at 100.111Code: https://github.com/dubai03nsr/dinco
@@ 1 Introduction
LLMs encode a vast amount of knowledge in their parameters, demonstrating superhuman performance on knowledge-intensive benchmarks (Comanici et al., 2025; OpenAI, 2023). Users often rely on information obtained from these models to make important decisions, but the information is not always accurate. Thus, we seek to qualify LLM responses with confidence estimates that are calibrated, i.e. match the probability of correctness. Users and agentic frameworks often use LLMs in a zero-shot manner without task-specific tuning (Manakul et al., 2023; Geng et al., 2024; Feng et al., 2024; Shorinwa et al., 2025), motivating the development of confidence estimation methods that work in off-the-shelf settings – both gray-box settings with logit access, and black-box settings with only textual input and output.
Figure 1: Calibration metrics (Expected Calibration Error $\downarrow$, Brier score $\downarrow$, area under the ROC curve $\uparrow$; see Appendix B.1) with Qwen3-8B on TriviaQA using $\rm P(True)$} as verbalized confidence. (Left) Verbalized confidence is saturated at high confidence. For each bar, we label the number of instances whose confidence falls in the interval and we darken larger bins. (Center) DiNCo normalizes by the total confidence over candidate answers, relieving saturation and improving calibration. (Right) Since verbalized confidence is saturated at high confidence, it is unable to achieve an acceptable true positive rate (TPR) without incurring a significant false positive rate (FPR) of 0.24. In other words, no rejection threshold can be chosen to reject a high proportion of false claims. Meanwhile, DiNCo enjoys better granularity, ranking positives above negatives even among instances with a verbalized confidence of 1.
In these settings, verbalized confidence is a simple and commonly-used approach that prompts the model to report its confidence in an answer (Lin et al., 2022; Xiong et al., 2024; Wei et al., 2024). For brevity, we use verbalized confidence as a blanket term for (1) asking the model to decode a numerical confidence like “80%” (Tian et al., 2023) and (2) asking the model whether an answer is correct and taking $\rm P(True)$} (Kadavath et al., 2022). Verbalized confidence is appealing for several reasons, including that it resembles one way humans express confidence, making it easy to interpret and integrate into decision-theoretic frameworks (Sun et al., 2025; Steyvers et al., 2025). However, verbalized confidence has several drawbacks. First, it empirically tends to exhibit overconfidence (Tian et al., 2023; Xiong et al., 2024; Wei et al., 2024; Xu et al., 2025); Fig. 1 (left) shows that verbalized confidence scores generally outstrip average accuracy within a confidence bin.
@@
To explain and correct for this kind of incoherence, we first define the notion of suggestibility. Some studies indicate that when LLMs are epistemically uncertain, they tend to rely on their context to resolve the uncertainty (Yadkori et al., 2024; Ahdritz et al., 2024), i.e., the confidence on a claim increases because it is in the context. We refer to this phenomenon as suggestibility and hypothesize that it contributes to the model’s assignment of high confidence to claims it can neither support nor refute. To account for this bias, we propose a method for calibrating verbalized confidence that normalizes by the total confidence over self-generated distractors (i.e. alternative claims). We generate minimal pair distractors using beam search when available, or by directly prompting the model for distractors in the black-box setting. Crucially, we use an off-the-shelf NLI model to downweight distractors that are similar to other distractors or that do not contradict the main claim.
@@
We test our method on open-source and closed-source models, applied to short-form (TriviaQA and SimpleQA) (Joshi et al., 2017; Wei et al., 2024) and long-form (FActScore; Min et al., 2023) generation domains. DiNCo improves ECE over the next best method by an average of 0.099, 0.092, and 0.055, respectively. DiNCo effectively extends to long-form biography generation, where it improves Pearson and Spearman correlation with passage-level FActScore over the best baseline by an average of 0.072 and 0.074, respectively. Further analysis shows that DiNCo relieves confidence saturation, and that simply scaling up self-consistency (the strongest baseline overall) does not suffice to match the calibration of DiNCo.
@@ 2.1 Background and Motivation
Figure 2: Normalizing verbalized confidence with DiNCo. (Left) The LLM generates a claim along with several distractors and reports its confidences on them independently. To calibrate the main claim’s confidence, we divide it by $\beta$, the sum over each distractor’s confidence, weighted by uniqueness (center) and counterfactuality (right). Details in Section 2.3.
@@ 2.2 Preliminary Study
To empirically support our motivation above, we plot the distributions of the total confidence ($\beta(C)$ in Eq. 2) over correctly and incorrectly answered questions. In this preliminary study, we take correctness as a proxy for epistemic certainty, although in reality LLMs may still be uncertain about questions they answered correctly. In other words, we treat correct and incorrect instances as epistemically certain and uncertain instances, respectively. Thus, our hypothesis predicts that on incorrect instances, the model would be more suggestible and assign high confidence to more answers, leading to higher total confidences. On the other hand, if the model is calibrated (in particular, on uncertain instances it exhibits epistemic humility, i.e. recognizing its lack of knowledge), then the total confidence would tend to be around 1 and 0 for correct and incorrect instances, respectively, so the total confidence would be lower on incorrect instances. Even if the model is confident in its incorrect answers (e.g. due to misconceptions), we would expect similar behavior between correct and incorrect instances.
@@
Figure 3: Distributions of total confidence over correct and incorrect answers.
In Section 2.3, we specify how we generate distractors and account for answer redundancy.
Results.  In Fig. 3, we first observe that many incorrect answers are assigned a confidence near 0, suggesting epistemic humility. Even so, the incorrect distribution is heavy-tailed, resulting in a higher mean and median than the correct distribution. These results are consistent with our hypothesis that LLMs are more prone to accepting claims that they are epistemically uncertain about.
@@ 2.3 Method
The distractor set should contain enough plausible distractors to avoid underestimating the normalization factor ($\beta(C)$ in Eq. 2), while remaining small enough to be tractably computed. Thus, we frame the problem of choosing the optimal distractor set containing an original claim $c_{0}$ as maximizing the total acceptance probability $\sum_{c\in C}f^{\rm VC}(c)$ subject to a size constraint $|C|\leq K$; shortly we address relaxing the requirement of mutual exclusivity. Unfortunately, the validation probability $f^{\rm VC}(c)$ can only be elicited on a per-claim basis, leading to intractable sample complexity. Motivated by the intuition that an LLM tends to generate claims that it would find plausible during validation, as a proxy for the set of claims with high verbalized confidence, we use the set of claims with high generation probability (under an appropriate prompt, explained below).
To encourage mutual exclusivity among claims, we set up the claims to be minimal pairs (Fig. 2 left). For a given short-form question, we can simply sample many answers, but independent sampling is inefficient as it overrepresents high probability generations (Gekhman et al., 2025), limiting the number of unique distractors. Instead, we use beam search (Sutskever et al., 2014) when available to efficiently identify unique sequences that have high probability mass coverage. For API-access models, we use top token probabilities if available to implement a pseudo-beam search (see Appendix C.1 for details). Otherwise in the black-box setting, we directly prompt the model to generate a list of candidate answers (see Appendix B.2 for prompts). For long-form QA, we follow Min et al. (2023) in decomposing a long generation into claims. We separately prompt the model to generate one distractor for a given claim, and again use beam search to create multiple distractors.
@@
Intuitively, $w_{\rm unique}$ downweights a claim if it is entailed by other claims, and $w_{\rm contra}$ downweights a claim if it is not contradictory with the main claim. We now normalize the verbalized confidence of the original claim as
@@ 3.1 Experimental Setup
Short-form Datasets.  Short-form QA serves as a testbed for evaluating factuality as well as calibration because of its tractable evaluation and adjustable difficulty. The task is relevant in practice because it assesses models’ ability to respond to information-seeking users. TriviaQA contains trivia questions requiring diverse world knowledge (Joshi et al., 2017). SimpleQA similarly contains short, fact-seeking questions, curated with the criterion of challenging frontier models (Wei et al., 2024). We sample 1000 questions from each dataset. [^6] We use the validation split of the rc.nocontext subset for TriviaQA. On SimpleQA, Gemini-2.5-Flash produced a refusal error with no output on 86 questions, so we exclude them from all experiments. We use LLM-as-a-judge to evaluate binary correctness, following best practices for robust evaluation (Wei et al., 2024); in Appendix B.3 we confirm high human agreement.
@@
Models.  Since TriviaQA and SimpleQA are adequately challenging for smaller and larger models, respectively, we focus their evaluation accordingly. TriviaQA is largely solved by frontier models (Wei et al., 2024), and SimpleQA is too difficult for smaller models. [^7] https://www.kaggle.com/benchmarks/openai/simpleqa On TriviaQA, we use popular open-source models: Qwen3-8B and Qwen3-1.7B (Yang et al., 2025), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Gemma-3-4B-IT (Team et al., 2025). On SimpleQA, we use popular frontier models: GPT-4.1 (2025-04-14; OpenAI, 2025) and Gemini-2.5-Flash (Comanici et al., 2025). For SimpleQA evaluated on frontier models, we also consider the black-box setting where no logit access is assumed; here, rather than using pseudo-beam search to generate distractors, we prompt the model directly to generate diverse distractors. Moreover, we replace $\rm P(True)$} with verbalized numerical confidence to forgo logit access. For the long-form task of biography generation, we limit our evaluation to Qwen3-8B and Gemma-3-4B-IT, due to the cost of FActScore evaluation with GPT-4.1. We use the NLI model DeBERTa-v3-base-mnli-fever-anli (He et al., 2021) for our methods and self-consistency.
@@
Evaluation Metrics.  We evaluate Expected Calibration Error (ECE $\downarrow$; Naeini et al., 2015) with 10 bins, Brier score (BS $\downarrow$; Brier, 1950), and area under the ROC curve (AUC $\uparrow$; Hanley & McNeil, 1982). See Appendix B.1 for descriptions. For biography generation, we also evaluate Pearson and Spearman correlation between average claim-level confidence and passage-level FActScore (without length penalty), where the latter measures the proportion of claims that are correct.
@@
Baselines.  Following past work in zero-resource calibration, we compare against training-free methods that produce probabilities without post-hoc calibration (Tian et al., 2023; Xiong et al., 2024; Steyvers et al., 2025). We provide prompts for our methods and baselines in Appendix B.2. Verbalized confidence (VC; $\rm P(True)$} from Kadavath et al., 2022) asks the model whether its answer is correct and computes $P({\rm Yes})/(P({\rm Yes})+P({\rm No}))$. It is straightforward to replace $\rm P(True)$} with verbalized numerical confidence; in Appendix D.1 we show that the latter similarly benefits from our methods, showing robustness to the format of verbalized confidence. Top-$K$ prompting ($K$-VC; Verb. 1S from Tian et al., 2024) prompts the model to provide its top $K$ guesses along with verbalized numerical confidences. For our black-box setting, we use the candidate answers generated using the $K$-VC prompt as distractors, but we discard the verbalized confidences contained in the same generation, and instead separately collect verbalized numerical confidence on each distractor independently. Maximum sequence probability (MSP; Fadeeva et al., 2023) is the LLM’s probability of generating its answer. Self-consistency (SC; Xiong et al., 2024) samples several answers (we use temperature 1) and computes the proportion that match the main answer. Following Kuhn et al. (2023), we use an NLI model to determine semantic equivalence.5 For biography generation, we modify self-consistency to sample several biographies and measure entailment of each claim (Zhang et al., 2024a). SC-VC (Xiong et al., 2024) is SC weighted by verbalized confidence (we use $\rm P(True)$} following Taubenfeld et al., 2025).
@@
Inference Budget.  As our methods and several baseline methods ($K$-VC, SC, SC-VC) operate on a variable inference-time budget, we control this budget at $K=10$. For DiNCo, we use 5 samples for self-consistency and 5 distractors for normalized verbalized confidence.

### v2

# arXiv:2509.25532v2[cs.CL] 27 Feb 2026

@@ Abstract
Calibrated confidence estimates are necessary for large language model (LLM) outputs to be trusted by human users. While LLMs can express their confidence in human-interpretable ways, verbalized LLM-generated confidence scores have empirically been found to be miscalibrated, reporting high confidence on instances with low accuracy and thereby harming trust and safety. We hypothesize that this overconfidence often stems from a given LLM’s heightened suggestibility when faced with claims that it encodes little information about; we empirically validate this hypothesis, finding more suggestibility on lower-accuracy claims. Building on this finding, we introduce Distractor-Normalized Coherence (DiNCo), which estimates and accounts for an LLM’s suggestibility bias by having the model verbalize its confidence independently across several self-generated distractors (i.e. alternative claims), and normalizes by the total verbalized confidence. To further improve calibration, we leverage generator-validator disagreement, augmenting normalized validator confidence with a consistency-based estimate of generator confidence. Here, we frame the popular approach of self-consistency as leveraging coherence across sampled generations, and normalized verbalized confidence as leveraging coherence across validations on incompatible claims, allowing us to integrate these complementary dimensions of coherence into DiNCo. Moreover, our analysis shows that DiNCo provides less saturated – and therefore more usable – confidence estimates, and that further sampling alone cannot close the gap between DiNCo and baselines, with DiNCo at 10 inference calls outperforming self-consistency at 100.111Code: https://github.com/victorwang37/dinco
@@ 1 Introduction
LLMs encode a vast amount of knowledge in their parameters, demonstrating superhuman performance on knowledge-intensive benchmarks (Comanici et al., 2025; OpenAI, 2023). Users often rely on information obtained from these models to make important decisions, but this information is not always accurate. Thus, we seek to qualify LLM responses with confidence estimates that are calibrated, i.e. match the probability of correctness. Users and agentic frameworks often use LLMs off the shelf without task-specific or model-specific tuning (Manakul et al., 2023; Geng et al., 2024; Feng et al., 2024; Shorinwa et al., 2025), motivating the development of confidence estimation methods that work in off-the-shelf settings – both gray-box settings with logit access, and black-box settings with only textual input and output.
Figure 1: Calibration metrics (Expected Calibration Error $\downarrow$, Brier score $\downarrow$, area under the ROC curve $\uparrow$; see Appendix B.1) with Qwen3-8B on TriviaQA using $\rm P(True)$} as verbalized confidence. (Left) Verbalized confidence is saturated at high confidence, despite a much lower accuracy. For each bar, we label the number of instances whose confidence falls in the interval and we darken larger bins. (Center) DiNCo normalizes by the total confidence over candidate answers, relieving saturation and improving calibration. (Right) Since verbalized confidence is saturated at high confidence, it is unable to achieve an acceptable true positive rate (TPR) without incurring a significant false positive rate (FPR) of 0.24. In other words, no rejection threshold can be chosen to reject a high proportion of false claims. Meanwhile, DiNCo enjoys better granularity, ranking positives above negatives even among instances with a verbalized confidence of 1.
In these settings, verbalized confidence is a simple and commonly-used approach that prompts the model to report its confidence in an answer (Lin et al., 2022; Xiong et al., 2024; Wei et al., 2024). For brevity, we use verbalized confidence as a blanket term for (1) asking the model to decode a numerical confidence in text like “80%” (Tian et al., 2023) and (2) asking the model whether an answer is correct and taking the token probability $\rm P(True)$} (Kadavath et al., 2022). Verbalized confidence is appealing for several reasons, including that it resembles one way humans express confidence, making it easy to interpret and integrate into decision-theoretic frameworks (Sun et al., 2025; Steyvers et al., 2025). However, verbalized confidence has several drawbacks. First, it empirically tends to exhibit overconfidence (Tian et al., 2023; Xiong et al., 2024; Wei et al., 2024; Xu et al., 2025); Fig. 1 (left) shows that verbalized confidence scores generally outstrip average accuracy within a confidence bin.
@@
To explain and correct for this kind of incoherence, we first define the notion of suggestibility. Some studies indicate that when LLMs are epistemically uncertain, they tend to rely on their context to resolve the uncertainty (Yadkori et al., 2024; Ahdritz et al., 2024), i.e., the confidence on a claim increases because it is in the context. We refer to this phenomenon as suggestibility and hypothesize that it contributes to the model’s assignment of high confidence to claims it can neither support nor refute. We introduce this hypothesis in Section 2.1 and provide empirical support in Section 2.2. To account for this suggestibility bias, we propose a method for calibrating verbalized confidence that normalizes by the total confidence over self-generated distractors (i.e. alternative claims). We generate minimal pair distractors using beam search when available, or by directly prompting the model for distractors in the black-box setting. Crucially, we use an off-the-shelf NLI model to downweight distractors that are similar to other distractors or that do not contradict the main claim.
@@
We test our method on open-source and closed-source models, applied to short-form (TriviaQA and SimpleQA) (Joshi et al., 2017; Wei et al., 2024) and long-form (FactScore; Min et al., 2023) generation domains. DiNCo improves ECE over the best baseline by an average of 0.077, 0.092, and 0.055, respectively (note that the best baseline differs between the short-form and long-form settings). DiNCo effectively extends to long-form biography generation, where it improves Pearson and Spearman correlation with passage-level FactScore over the best baseline by an average of 0.072 and 0.074, respectively. Further analysis shows that DiNCo relieves confidence saturation, and that simply scaling up self-consistency (the strongest baseline overall) does not suffice to match the calibration of DiNCo.
@@ 2.1 Background and Motivation
Figure 2: Normalizing verbalized confidence with DiNCo. (Left) The LLM generates a claim along with several distractors and reports its confidences on them independently. To calibrate the main claim’s confidence, we divide it by $\beta$, the sum over each distractor’s confidence, weighted by uniqueness (center) and counterfactuality (right). We seek distractors that are minimal pairs with the main claim, i.e. similar statements that likely contradict the main claim. The distractors shown here are real examples generated by an LLM, and while they happen to share a lexical structure with the main claim, we do not prescribe a precise form for generating distractors; see Section 2.3 for method details, Appendix B.2 for prompts, and Appendix C.1 for more examples and analysis.
@@ 2.2 Preliminary Study
To empirically support our motivation above, we plot the distributions of the total confidence ($\beta(C)$ in Eq. 2) over correctly and incorrectly answered questions. In this preliminary study, we take correctness as a proxy for epistemic certainty, although in reality LLMs may still be uncertain about questions they answered correctly. In other words, we treat correct and incorrect instances as epistemically certain and uncertain instances, respectively. Thus, our hypothesis predicts that on incorrect instances, the model would be more suggestible and assign high confidence to more answers, leading to higher total confidences. On the other hand, if the model is calibrated (in particular, on uncertain instances, it exhibits epistemic humility, i.e. recognizing its lack of knowledge), then the total confidence would tend to be around 1 and 0 for correct and incorrect instances, respectively, so the total confidence would be lower on incorrect instances. If the model is confident in its incorrect answers (e.g. due to misconceptions), we would expect similar behavior between correct and incorrect instances.
@@
Figure 3: Total confidence for correctly and incorrectly answered questions.
For each question, we generate 10 claims, so the total confidence is a number from 0 to 10. In Section 2.3, we specify how the distractors are generated and how we avoid overcounting redundant answers in the total confidence (Eq. 4). The total confidence $\beta(C)$ computed here matches how the NVC method and the NVC component of DiNCo (but with fewer distractors) will be computed for our main experiments in Section 3.
Results.  In Fig. 3, we first observe that some incorrectly answered questions have a total confidence near 0, suggesting some epistemic humility. Even so, the incorrect distribution is heavy-tailed, resulting in a higher mean and median than the correct distribution. These results are consistent with our hypothesis that LLMs are more prone to accepting claims that they are epistemically uncertain about. To explain why the model experiences suggestibility even on correct instances (shown by the substantial proportion of correct instances with a total confidence greater than 1), we note that educated guesses can often be correct while still bearing epistemic uncertainty. Despite our usage of correctness as a noisy proxy of epistemic certainty, we identify the clear trend that the model tends to report higher confidence on claims for incorrectly answered questions. We verify the same trend on SimpleQA (Wei et al., 2024) and biography generation (Min et al., 2023) in Appendix D.1.
@@ 2.3 Method
The distractor set should contain enough plausible distractors to avoid underestimating the normalization factor ($\beta(C)$ in Eq. 2), while remaining small enough to be tractably computed. Thus, we frame the problem of choosing the optimal distractor set containing an original claim $c_{0}$ as maximizing the total acceptance probability $\sum_{c\in C}f^{\rm VC}(c)$ subject to a size constraint $|C|\leq K$; shortly we address relaxing the requirement of mutual exclusivity. Unfortunately, the validation probability $f^{\rm VC}(c)$ can only be elicited on a per-claim basis, leading to intractable sample complexity. Motivated by the intuition that an LLM tends to generate claims that it would find plausible during validation, as a proxy for the set of claims with high verbalized confidence, we use the set of claims with high generation probability (under an appropriate prompt, explained next).
To encourage mutual exclusivity among claims, we set up the claims to be minimal pairs (Fig. 2 left). For a given short-form question, we can simply sample many answers, but independent sampling is inefficient as it overrepresents high probability generations (Gekhman et al., 2025), limiting the number of unique distractors. Instead, we use beam search (Sutskever et al., 2014) when available to efficiently identify unique sequences that have high probability mass coverage. For API-access models, we use top token probabilities if available to implement a pseudo-beam search (see Appendix C.2 for details). Otherwise in the black-box setting, we directly prompt the model to generate a list of candidate answers (see Appendix B.2 for prompts). For long-form QA, we follow Min et al. (2023) in decomposing a long generation into claims. We separately prompt the model to generate one distractor for a given claim, and again use beam search to create multiple distractors. We provide examples and analysis of the distractors generated with each of these methods in Appendix C.1.
@@
Intuitively, $w_{\rm unique}$ downweights a claim if it is entailed by other claims, and $w_{\rm contra}$ downweights a claim if it is not contradictory with the main claim. In the simplified setting where claims are either fully equivalent or fully contradictory, $w_{\rm unique}$ for a claim is the reciprocal of the size of its equivalence class, so we have invariance to claim duplication. Similarly, $w_{\rm contra}$ grants invariance to including the original claim as a distractor. Beyond this simplified setting, using continuous weights allows us to model partial entailment or contradiction, such as for the claims “Kang Ji-hwan is widely ridiculed.” and “Kang Ji-hwan is notoriously disliked.” in Fig. 2.
We now normalize the verbalized confidence of the original claim as
@@ 3.1 Experimental Setup
Short-form Datasets.  Short-form QA serves as a testbed for evaluating factuality as well as calibration because of its tractable evaluation and adjustable difficulty. The task is relevant in practice because it assesses models’ ability to respond to information-seeking users. TriviaQA contains trivia questions requiring diverse world knowledge (Joshi et al., 2017). SimpleQA similarly contains short, fact-seeking questions, curated with the criterion of challenging frontier models (Wei et al., 2024). We sample 1000 questions from each dataset. [^6] We use the validation split of the rc.nocontext subset for TriviaQA. On SimpleQA, Gemini-2.5-Flash produced a refusal error with no output on 86 questions, so we exclude them from all experiments. In Appendix D.6, we evaluate on BioASQ (Krithara et al., 2023), a dataset demanding biomedical expertise, to show generalization to other domains. We use LLM-as-a-judge to evaluate binary correctness, following best practices for robust evaluation (Wei et al., 2024); in Appendix B.3 we confirm high human agreement.
@@
Models.  Since TriviaQA and SimpleQA are adequately challenging for smaller and larger models, respectively, we focus their evaluation accordingly. TriviaQA is largely solved by frontier models (Wei et al., 2024), and SimpleQA is too difficult for smaller models. [^7] https://www.kaggle.com/benchmarks/openai/simpleqa On TriviaQA, we use popular open-source models: Qwen3-32B, Qwen3-8B, and Qwen3-1.7B (Yang et al., 2025), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Gemma-3-4B-IT (Team et al., 2025). On SimpleQA, we use popular frontier models: GPT-4.1 (2025-04-14; OpenAI, 2025) and Gemini-2.5-Flash (Comanici et al., 2025). For SimpleQA evaluated on frontier models, we also consider the black-box setting where no logit access is assumed; here, rather than using pseudo-beam search to generate distractors, we prompt the model directly to generate diverse distractors. Moreover, we replace $\rm P(True)$} with verbalized numerical confidence to forgo logit access. For the long-form task of biography generation, we limit our evaluation to Qwen3-8B and Gemma-3-4B-IT, due to the cost of FactScore evaluation with GPT-4.1. We use the NLI model DeBERTa-v3-base-mnli-fever-anli (He et al., 2021) for our methods and self-consistency; we confirm robustness to the choice of the NLI model in Appendix D.3.
@@
Evaluation Metrics.  We evaluate Expected Calibration Error (ECE $\downarrow$; Naeini et al., 2015) with 10 bins, Brier score (BS $\downarrow$; Brier, 1950), and area under the ROC curve (AUC $\uparrow$; Hanley & McNeil, 1982). See Appendix B.1 for descriptions. Like accuracy, these calibration metrics are on a scale from 0 to 1, meaning that e.g., an improvement of 0.05 is substantial, as it corresponds to a 5% absolute improvement (Tian et al., 2023; Xiong et al., 2024); we report all improvements in absolute rather than relative terms. For biography generation, we also evaluate Pearson and Spearman correlation between average claim-level confidence and passage-level FactScore (without length penalty), where the latter measures the proportion of claims that are correct.
@@
Baselines.  Following past work in zero-resource calibration, we compare against training-free methods that produce probabilities without post-hoc calibration (Tian et al., 2023; Xiong et al., 2024; Steyvers et al., 2025). We provide prompts for our methods and baselines in Appendix B.2. Verbalized confidence (VC; $\rm P(True)$} from Kadavath et al., 2022) asks the model whether its answer is correct and computes $P({\rm Yes})/(P({\rm Yes})+P({\rm No}))$. It is straightforward to replace $\rm P(True)$} with verbalized numerical confidence; in Appendix D.2 we show that the latter similarly benefits from our methods, showing robustness to the format of verbalized confidence. Top-$K$ prompting ($K$-VC; Verb. 1S from Tian et al., 2023) prompts the model to provide its top $K$ guesses along with verbalized numerical confidences. For our black-box setting, we use the candidate answers generated using the $K$-VC prompt as distractors, but we discard the verbalized confidences contained in the same generation, and instead separately collect verbalized numerical confidence on each distractor independently. Maximum sequence probability (MSP; Fadeeva et al., 2023) is the LLM’s probability of generating its answer. Self-consistency (SC; Xiong et al., 2024) samples several answers (we use temperature 1) and computes the proportion that match the main answer. Following Kuhn et al. (2023), we use an NLI model to determine semantic equivalence when grouping together answers for SC.5 For biography generation, we modify self-consistency to sample several biographies and measure entailment of each claim (Zhang et al., 2024a). SC-VC (Xiong et al., 2024) is SC weighted by verbalized confidence (we use $\rm P(True)$} following Taubenfeld et al., 2025).
@@
Inference Budget.  As our methods and several baselines ($K$-VC, SC, SC-VC) operate on a variable inference-time budget, we control this budget at $K=10$. For DiNCo, we use 5 samples for self-consistency and 5 distractors for normalized verbalized confidence; we show robustness to the budget split for various budgets in Appendix D.4 and thus recommend an equal split for simplicity.